# Domain Generalization in-the-Wild: Disentangling Classification from Domain-Aware Representations

### v1

@@ 6 Conclusion
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and/or introduction clearly states the claims made, including the contributions made in the paper and important assumptions and limitations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations mentioned in separate section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No proofs included.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information is disclosed.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Data and code will be public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All information is provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Statistical significance is provided for some results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Hardware information is mentioned.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Only publicly available datasets and models were used.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Societal impact not mentioned.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
[…]

### v2

@@ 5 Conclusion
Appendix A Extended Related Work
Domain Generalization. Learning domain-invariant representations has historically been a central idea in domain generalization (Blanchard et al., 2011; Zhou et al., 2022a). The intuition is that when classifying images from entirely new distributions, learning abstract features common across source domains should provide better robustness for classification in new domains (Blanchard et al., 2011; Muandet et al., 2013). Among these, domain-adversarial learning methods have become a relatively standard approach within the DG field due to its conceptual simplicity and effectiveness (Zhou et al., 2022a). For instance, Domain Adversarial Neural Networks (DANN) (Ganin et al., 2016) uses an auxiliary domain classifier trained adversarially against the encoder, encouraging the encoder to produce features indistinguishable across source domains. Given the focus of DANN on the central idea of domain invariance, we focus on DANN and its adaptation to CLIP in our analysis. Notably, despite the prevalence of such DG methods, the direct application for CLIP is not well-established and remains underexplored. Naively enforcing domain invariance on foundation models like CLIP, with large pretrained knowledge, risks catastrophic forgetting.
Robust Finetuning of CLIP. The introduction of CLIP marked a significant shift in DG research. The original study (Radford et al., 2021) demonstrated impressive zero-shot classification performance across diverse benchmarks, including OOD datasets. The authors attributed this capability to CLIP learning representations that are less reliant on spurious correlations specific to downstream target datasets, as CLIP was not trained on these specific datasets during its initial pretraining.
The assumption of the inherent OOD robustness in CLIP motivated numerous methods aimed at finetuning CLIP for downstream tasks while enhancing its perceived robustness. A common approach is parameter-efficient finetuning (PEFT) strategies. An early influential study, CoOp (Zhou et al., 2022c), introduced learnable textual prompts, motivated by observations that manually crafted prompt ensembles improved CLIP’s zero-shot accuracy. Building on this, CoCoOp (Zhou et al., 2022b) made these prompts dynamic by conditioning them on individual image features through a cross-attention mechanism. Similarly, CLIP-Adapter (Gao et al., 2024) proposed adding lightweight, learnable MLP layers (adapters) to the CLIP encoders, finetuning only these small adapters instead of the entire network. Many more subsequent PEFT methods have also been explored (Cho et al., 2023; Chi et al., 2024; Lee et al., 2025; Addepalli et al., 2024; Bai et al., 2024; Li et al., 2022; Khattak et al., 2025; Cheng et al., 2024; Lafon et al., 2024).
End-to-end finetuning methods have also been explored, yet many still depend on the original pretrained CLIP weights for regularization or guidance. Wise-FT (Wortsman et al., 2022), motivated by observing that standard finetuning often degraded zero-shot OOD performance, ensembles the weights of the finetuned model with the original CLIP weights. CLIP-OOD (Shu et al., 2023) used a beta-moving average of the weights during finetuning alongside a regularization term to enhance semantic relationships learned during pretraining. MIRO (Cha et al., 2022) used mutual information regularization between the finetuning model and the frozen pretrained CLIP model to retain pretrained features.
While many other methods show strong performance on OOD benchmarks, this overview highlights representative approaches, their trends, and assumptions in robust CLIP finetuning. Our work, however, questions whether current evaluation protocols are sufficiently challenging, and suggests the reliance on the pretrained weights may be suboptimal for true OOD generalization, a concern supported by evidence of domain contamination during pretraining (Mayilvahanan et al., 2024). Consequently, we explore more challenging evaluations and alternative strategies for training CLIP grounded in DG principles.
Appendix B Pseudocode
B.1 Main Training Loop (CLIP-DCA)
1# Note: logit scales (temperature term) were omitted for simplicity
2def classification_loss(image, text):
3 logits_per_image = image @ text.T
4 logits_per_text = text @ image.T
5 labels = torch.arange(len(image), device=device)
6 return (F.cross_entropy(logits_per_image, labels) +
7 F.cross_entropy(logits_per_text, labels)) / 2
8
9def disentangle_loss(x, y):
10 x = (x - x.mean(0)) / (x.std(0) + 1e-8)
11 y = (y - y.mean(0)) / (y.std(0) + 1e-8)
12 cross_cor_mat = (x @ y.T) / len(x)
13 return torch.diagonal(cross_cor_mat).pow(2).sum()
14
15# Architectural additions (training only)
16domain_head = $n$.Linear(...)
17mllm_projector = $n$.Linear(...)
18
19# 1. Diffusion data batch
20d_images, d_hidden, d_text = d_batch
21penultimate_img_emb, class_img_emb, text_emb = clip_model(d_images, d_text)
22domain_img_emb = domain_head(penultimate_img_emb.clone())
23domain_mllm_emb = mllm_projector(d_hidden)
24
25C5 = classification_loss(domain_img_emb, text_emb) # Domain Agreement
26C6 = classification_loss(domain_mllm_emb, text_emb) # MLLM Agreement
27C3 = disentangle_loss(class_img_emb, domain_img_emb)
28C4 = disentangle_loss(class_img_emb, text_emb)
29
30loss_diffusion = C3 + C4 + C5 + C6
31
32# 2. Source data batch
33s_images, s_text = s_batch
34penultimate_img_emb, class_img_emb, text_emb = clip_model(s_images, s_text)
35domain_img_emb = domain_head(penultimate_img_emb.clone())
36
37C1 = classification_loss(class_img_emb, text_emb) # Classification Loss
38C2 = disentangle_loss(class_img_emb, domain_img_emb)
39
40loss_source = C1 + C2
41
42# 3. Final loss
43loss = loss_diffusion + loss_source
Listing 1: Main training loop for CLIP-DCA.
B.2 Unlearning Loop
1# Note: logit scales (temperature term) were omitted for simplicity
2discriminator = $n$.Linear(...)
3
4# Following Ganin et al. (2016)
5p = float(batch_idx + start_steps) / total_steps
6alpha = 2. / (1. + np.exp(-10 * p)) - 1
7
8def classification_loss(image, text):
9 # ... (same as in Algorithm 1)
10
11# 1. Retention on GCC dataset
12gcc_images, gcc_captions = gcc_batch
13_, image_emb, text_emb = clip_model(gcc_images, gcc_captions)
14retention_loss = classification_loss(image_emb, text_emb)
15
16# 2. Unlearning on DomainNet
17dn_images, _ = dn_batch
18noise = torch.randn_like(dn_images, requires_grad=True)
19
20# Pass both through encoder; apply gradient reversal to discriminator
21features_dn, _ = clip_model.encode_image(dn_images)
22features_noise, _ = clip_model.encode_image(noise)
23
24combined_features = torch.cat((features_dn, features_noise), dim=0)
25reversed_features = GradientReversalFunction.apply(combined_features, alpha)
26
27domain_logits = discriminator(reversed_features)
28domain_targets = torch.cat((torch.ones(len(dn_images)),
29 torch.zeros(len(noise))), dim=0).long()
30unlearning_loss = F.cross_entropy(domain_logits, domain_targets)
31
32# 3. Final loss
33loss = retention_loss + unlearning_loss
Listing 2: Unlearning loop with gradient reversal.
Appendix C Training Details
For all experiments, we used the CLIP ViT-B/32 model. Models were finetuned on the ImageNet-1k training set for 5 epochs. The official ImageNet validation set was used. We used the AdamW optimizer with a learning rate of 1e-5, with a cosine learning rate scheduler. Due to computational constraints, a consistent batch size of 128 was maintained across all methods. For baseline methods, any additional method-specific hyperparameters were adopted from the default configurations provided in their publicly available codebases. All experiments were conducted with a single NVIDIA RTX A5000 GPU and an AMD EPYC 7763 CPU.
Appendix D Hyperparameter tuning
This section provides details for hyperparameter tuning. In the main manuscript, we report 6 different losses for the distinction between source data and diffusion generated data ($C_{1}$ - $C_{6}$). For hyperparameter tuning, we group these losses into four terms. $C_{2}$ and $C_{3}$ are grouped together as the image disentangle term, while $C_{1}$ and $C_{4}$ are grouped together as the text disentangle term. We report our hyperparameter tuning in Table . Due to computation limitation, the tuning was limited to single term increments.
Table D.1: Average accuracy across 33 datasets for specific hyperparameter combinations.
MLLM hidden state MLLM description Text disentangle Image disentangle Avg. Acc. (%)
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $50.8$
$\text{\times}{10}^{-3} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $49.6$
$\text{\times}{10}^{-2} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $49.0$
$\text{\times}{10}^{-1} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $50.2$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-3}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $51.3$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-2}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $50.4$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-1}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $49.6$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-3} 1×1$$\text{\times}{10}^{-4}$ $52.1$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-2} 1×1$$\text{\times}{10}^{-4}$ $51.4$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-1} 1×1$$\text{\times}{10}^{-4}$ $51.0$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-3}$ $50.6$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-2}$ $49.6$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-1}$ $49.1$
$\text{\times}{10}^{-4} 1×1$$\text{\times}{10}^{-4}$ $\text{\times}{10}^{-3} 1×1$$\text{\times}{10}^{-4}$ $52.1$ (Best)
Appendix E Datasets
We report all target datasets used for our experiments. Note that for training, the ImageNet-1K training set was used.
Table E.1: List of datasets used in experiments, including number of classes and images. DG benchmarks are listed first. Each domain within a DG benchmark is treated as a distinct dataset. The number of images represent the validation/test set.
Dataset Name Brief Description # Classes # Images
Domain Generalization (DG) Benchmarks
Digits DG Collection of digit recognition datasets.
MNIST Grayscale handwritten digits (28x28). 10 6,000
MNIST-M MNIST digits with color patches blended from BSDS500. 10 6,000
SVHN Colored house numbers from Google Street View images (32x32). 10 6,000
SYN Synthetically generated digit images (32x32). 10 6,000
Terra Incognita Camera trap images of wild animals from different locations.
Location 100 Animal images from Location 100. 10 4,741
Location 38 Animal images from Location 38. 10 9,736
Location 43 Animal images from Location 43. 10 3,970
Location 46 Animal images from Location 46. 10 5,883
PACS Object recognition with domain shifts.
Art Painting Artistic paintings of objects. 7 2,048
Cartoon Cartoon images of objects. 7 2,344
Photo Photographic images of objects. 7 1,670
Sketch Sketch drawings of objects. 7 3,929
Office-Home Object recognition in different settings.
Art Artistic depictions of everyday objects. 65 1,972
Clipart Clipart images of everyday objects. 65 3,910
Product Product images of everyday objects (typically clean backgrounds). 65 3,984
Real Real-world photographic images of everyday objects. 65 3,902
Individual Benchmark Datasets
Caltech-101 101 object categories (+1 background). 101 8,677
Oxford-IIIT Pets Images of pet breeds. 37 3,669
Oxford Flowers 102 Images of flower categories. 102 6,149
Stanford Cars Images of car makes, models, and years. 196 8,041
Food-101 Images of food categories. 101 25,250
FGVC Aircraft Images of aircraft variants. 100 3,333
SUN397 Scene understanding dataset with scene categories. 397 108,754
Describable Textures Dataset (DTD) Textures in the wild, organized by 47 human-perceivable attributes. 47 1,880
EuroSAT Satellite imagery of land use and land cover classes. 10 27,000
UCF101 Action recognition dataset of human action categories from videos. 101 13,320
ImageNet-1K 1.28M natural images in 1000 classes (ILSVRC 2012). 1000 50,000
ImageNet-V2 New test set for ImageNet-1K. 1000 50,889
ImageNet-Sketch Sketch images corresponding to ImageNet-1K. 1000 50,000
ImageNet-A "Natural adversarial examples" of 200 classes. 200 7,500
ImageNet-R "Renditions" (art, cartoons, etc.) of 200 ImageNet classes. 200 30,000
WILDS Benchmark Datasets (Treated as Individual)
Camelyon17-Wilds Histopathological images for tumor detection with hospital-based shifts. 2 85,054
FMOW-Wilds Satellite imagery for land use classification with temporal/regional shifts. 62 53,473
Appendix F Prompts Used for Multi-Modal Language Models (MLLMs)
This section details the specific prompts provided to Multi-Modal Language Models (MLLMs) for the generation tasks.
F.1 Prompt to MLLM to generate ideas for different styles of images
The following prompt was used to instruct the MLLM to generate a diverse list of image style ideas:
Give me ideas of 512 different styles of images.
Each style should be less than 5 words. Do not overlap styles.
Make the styles diverse.
Be brief.
F.2 Prompt to MLLM to generate descriptions and hidden states
The following prompt was used to instruct the MLLM to generate detailed descriptions of image styles (independent of object category) and to also extract corresponding hidden states. The following prompt was input together with images in each style:
Attached are multiple images in the same style.
Describe the aspects of the style that applies regardless of category.
Provide a description.
Do not describe the object in the image, but the style of image.
Be as detailed, complete, and comprehensive as possible.
Explain every minute detail.
Appendix G List of all synthetic styles
This section provides the list of all synthetic style dieas that were generated by the LLM.
Table LABEL:tab:synthetic_styles_list shows the 512 distinct style prompts used for generating synthetic data. The styles are listed alphabetically across the columns.
Table G.1: List of 512 synthetic data generation styles (alphabetical order).
3D rendered image 3D rendering, virtual objects ASCII art text ASCII art, text characters
Aboriginal dot painting Abstract expressionism Abstract expressionism art Abstract symbolic representation
Abstract, non-representational Abstract, non-representational form Achromatic grayscale image Achromatic, no color
Acrylic paint vibrant Action painting, dynamic Aerial drone footage Afrofuturism cultural sci-fi
Algorithmic art, code-based Ambient light, natural tones American scene painting Analog film, imperfections
Analogous colors harmony Anamorphic distorted perspective Ancient Egyptian hieroglyphs Animal at rest
Animal drinking, water source Animal eye contact Animal fighting, intense conflict Animal grooming, self-care
Animal hiding, partially obscured Animal hunting, focused gaze Animal looking away Animal marking territory
Animal mid-stride Animal playing, energetic Animal sleeping, peaceful Animal tracks, foreground focus
Animal vocalizing, mouth open Anime Japanese animation Anime, Japanese animation Architectural, building structures
Art Deco geometry Art Nouveau curves Artificial light controlled Arts and Crafts
Assemblage found objects Assemblage, 3D collage Astrophotography star trails Asymmetrical dynamic balance
Asymmetry, unbalanced design Augmented reality, overlaid Autumn leaves, warm palette Available light natural
Avant-garde, experimental Backlit silhouette lighting Backlit subject, glowing outline Baroque dramatic lighting
Bio art, living organisms Biopunk organic technology Biopunk, genetic engineering Bird’s-eye view elevated
Bird’s-eye view, distant Black and white film Blacklight fluorescent colors Blooming flowers, vibrant colors
Blue hour twilight Blueprint architectural plan Body art, human canvas Bokeh light effect
Bold geometric patterns Boomerang, looping video Botanical art, plant subjects Bright cheerful aesthetic
Broad lighting face Butterfly lighting beauty Byzantine mosaic icons Calligraphy elegant lettering
Calligraphy, elegant handwriting Camera flash, harsh light Camouflaged animal, hidden Candid street photography
Candid, unposed moment Caricature exaggerated features Cartoon simplified drawing Cartoon, exaggerated features
Cave entrance, dark frame Cave painting prehistoric Charcoal sketch drawing Charcoal sketch, rough lines
Chibi cute style Chromatic aberration, color fringing Cinemagraph, subtle movement Claymation stop-motion animation
Clear sky, bright blue Close up macro Close-up, animal portrait Close-up, feather detail
Close-up, fur texture Close-up, scale pattern Cloudscape art, sky scenes Collage mixed media
Collage, mixed media Color contrast, complementary Color field painting Color grading cinematic
Color splash accent Color temperature cool Color temperature warm Comic book style
Comic book, panel style Complementary colors contrast Conceptual, idea-driven Cross polarization, vibrant colors
Cross-hatching line work Cross-processed film Cross-processed, altered colors Cubism, geometric shapes
Cubist geometric forms Cyanotype process print Cyanotype, blue print Cyberpunk cityscape night
Cyberpunk style, dystopian future Daguerreotype antique look Dark ominous undertones Data bending corrupted
Data visualization, information art Decorative, ornamental style Decoupage, glued paper cutouts Deep focus sharp
Delicate fine details Dense jungle, lush green Depth of field Depth of field, blurred/sharp
Desaturated muted colors Desaturated, almost monochrome Dieselpunk retro-futuristic Different species together
Diffuse reflection matte Digital art, computer-generated Digital glitchy aesthetic Digital noise, grain effect
Digital painting software Digital print, inkjet/laser Distortion, warped perspective Documentary style, realistic
Doodle art casual Double exposure overlay Double exposure, ghost image Dramatic low-key lighting
Dramatic sky, storm clouds Dramatic spotlight, single source Dreamy ethereal soft focus Duotone color scheme
Duotone, two-color palette Dusty trail, arid environment Dutch angle tilted Dynamic energetic composition
Earth art, natural materials Embroidery thread texture Engraving detailed metal Engraving, incised lines
Environmental art, nature-focused Environmental portrait, surroundings Establishing shot context Etching fine lines
Etching, acid-etched lines Expressionist bold colors Extreme close-up detail Extreme close-up, eye detail
Fantasy art, mythical creatures Fashion illustration stylish Fashion, clothing focus Fast motion, sped-up action
Fauvist wild beasts Feeding animals, close action Film noir style Fine art, aesthetic focus
Fish-eye lens view Fish-eye lens, distorted Flat lay top-down Flowing river, blurred water
Focus stacking, all sharp Foggy morning, atmospheric haze Folk art naive Folk art, traditional craft
Forced perspective trick Forced perspective, size illusion Formal, posed shot Found object art
Fractal art mathematical Fractal art, mathematical Frontlit subject, clear view Frozen lake, icy surface
Full shot composition Futuristic sci-fi vision Futuristic style, advanced Generative art algorithmic
Generative art, algorithms Geometric abstract pattern Glamour, idealized beauty Glassblowing molten glass
Glitch art digital Glitch art distortion Glitch art, corrupted data Glitch art, digital errors
Golden hour sunlight Golden ratio composition Golden ratio, proportions Gothic art, dark, romantic
Gothic dark shadows Gouache opaque matte Graffiti art, street tagging Graffiti wildstyle lettering
Graffiti, tagged look Grainy film texture Grainy film, retro style Graphic design, visual communication
Gritty black and white Gritty urban decay Group of animals, social Group portraiture, multiple people
HDR photo rendering HDR, high dynamic range Halation, glowing highlights Hand-drawn sketchy feel
Hard light defined Hard light shadows Heavily textured impasto High contrast, dramatic lighting
High saturation, vivid colors High-angle shot looking High-key bright lighting High-speed photography
Holographic iridescent effect Horror art, scary imagery Hudson River School Hyperlapse, moving time-lapse
Hyperrealism, beyond realism Illuminated manuscript gold Illustrative, narrative imagery Impressionism, loose brushstrokes
Impressionist brushstrokes Industrial mechanical elements Infographic data visualization Infrared luminescence, glowing foliage
Infrared photography Infrared, false color Infrared, heat signature Ink wash fluid
Installation art, three-dimensional Interactive art, participation Isometric projection view Jewelry intricate design
Kinetic art movement Kinetic art, movement Land art earthworks Landscape art, natural scenery
Leading lines perspective Leading lines, guide eye Lens flare sunlight Lens flare, bright streaks
Light art, illumination Light leak, color streaks Light painting trails Line art contour
Line drawing, simple outline Linocut bold lines Lithograph stone print Lithography, planographic print
Lomography film look Long exposure shot Long exposure, light trails Long shot distance
Loop lighting portrait Low angle, animal towering Low poly geometric Low saturation, muted tones
Low-angle shot upwards Low-key dark lighting Lowbrow art, underground Lowbrow pop surrealism
Luminism glowing light Macro lens close-up Macro shot, tiny details Mandala circular symmetry
Manga graphic novel Map cartographic representation Maximalist busy composition Maximalist, elaborate design
Medium shot framing Metalwork shaped metal Migrating herd, vast landscape Miniature diorama world
Miniature effect, tilt-shift Minimalism, essential elements Minimalist simple lines Minimalist, simplified design
Mixed lighting combined Monochrome single color Monochrome, single color Monotype, unique print
Moody atmospheric lighting Moonlit night, stark shadows Mosaic tile pieces Mosaic, small piece patterns
Motion blur capture Motion blur, animal running Motion blur, speed lines Mountain range, panoramic view
Multiple exposure, layered images Mural art, large-scale painting Mural large-scale painting Naive art, childlike simplicity
Natural organic forms Negative space drawing Negative space framing Negative space, empty area
Neoclassical refined style Neon light glowing Nesting birds, detailed feathers Night photography cityscape
Night vision, green tint Night vision, red tint Oil painting texture Oil painting, thick texture
Op Art optical Op art, visual illusions Optical illusion, trickery Origami, paper folding
Ornate intricate design Orthochromatic film effect Outrun style neon Outsider art raw
Outsider art, untrained Over-the-shoulder perspective shot Overcast sky diffusion Overexposed, bright white
Overgrown, vegetation focus Panning motion blur Panning, blurred background Panoramic stitched view
Panoramic, wide environment Paper cut layered Paper marbling, swirling patterns Parent and offspring
Pastel drawing, blended colors Pastel soft blending Pattern repetition, visual rhythm Pen and ink
Pencil shading detailed Performance art, live action Photojournalism, storytelling Photorealism hyper-detailed
Photorealism, lifelike detail Pinhole camera image Pixel art retro Pixel art, retro game
Pixelated low resolution Pixelated, low resolution Point-of-view subjective shot Pointillism, tiny dots
Pointillist dot technique Polaroid transfer, image manipulation Polychrome, many colors Pop art bright
Pop art, bold colors Portraiture, individual likeness Positive space, subject focus Pottery ceramic art
Pre-Raphaelite detailed beauty Predator-prey interaction Psychedelic art, mind-altering Quilling, paper filigree
Quilting patchwork design Rack focus shift Radial balance, circular focus Rainy day, blurred drops
Realist everyday life Rembrandt lighting portrait Renaissance classical style Retro style, vintage look
Rocky terrain, jagged edges Rococo ornate details Romantic emotional landscape Rule of thirds
Rule of thirds, composition Rustic textured surface Sandy desert, dunes stretch Saturated vibrant colors
Schematic diagram layout Sci-fi art, space, technology Screen printing bold Screen printing, stencil print
Sculpture three-dimensional form Seascape art, ocean views Selective color isolation Selective focus, sharp animal
Self-portraiture, artist’s image Sepia tone, vintage look Sepia toned photograph Serene calming atmosphere
Shallow focus blur Sharp contrasting lines Sharp focus, crisp details Short exposure, frozen motion
Short lighting slimming Sidelit subject, defined features Silhouette backlit subject Silhouette black shape
Silhouette, dark shape Single-point lighting setup Sleek modern minimalist Slow motion, extended time
Smooth airbrushed finish Smooth digital, clean look Snowy scene, whiteout effect Social realism commentary
Soft focus, dreamy effect Soft light diffused Soft pastel hues Softbox diffused light
Solar punk green Solitary animal, minimalist Sound art, auditory Sparse woodland, bare trees
Specular highlights reflections Split lighting dramatic Split toning effect Split toning, colored highlights/shadows
Spring growth, fresh shoots Square color scheme Staged photography Stained glass colorful
Steampunk Victorian sci-fi Steampunk style, Victorian sci-fi Stencil art spray Stencil art, cut-out shapes
Stippling dot shading Stop motion, frame-by-frame Street art graffiti Street art, urban style
Studio portrait lighting Sumi-e ink painting Summer heat, shimmering air Sunlit glade, dappled light
Sunrise glow, warm tones Sunset silhouette, golden hour Surreal, bizarre, dreamlike Surrealism, dreamlike imagery
Surrealist dreamlike scene Symmetrical balanced framing Symmetry, balanced image Technical drawing precise
[…]